Circular No. 7017

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

V1333 AQUILAE S. A. Ilovaisky and C. Chevalier, Observatoire de Haute-Provence

(OHP), write: “CCD observations of the optical counterpart of the Aquila X-1 recurrent x-ray transient have obtained with the 1.2-m OHP telescope during Sept. 8 to 18. A V -band image taken on Sept. 10.86 UT shows the object to be still in quiescence, but on images taken by R. Mujica (Observatoire de Strasbourg) on Sept. 16.8 and 18.8, the object is 30 and 90 percent above the quiescent level, respectively. This brightening comes 200 days after the March episode (IAUC 6828, 6832), and it probably signals the start of a new activity cycle. Monitoring will continue at OHP. Observations at other wavelengths are encouraged.”

SUPERNOVA 1998bw IN ESO 184-G82 F. Patat, European Southern Observatory (ESO); and A. Piemonte,

Pontificia Universidad Catolica, on behalf of the ESO team (Lidman et al., cf. IAUC 6895, plus J. Brewer), report: “A low-resolution spectrum (range 350–1000 nm, resolution 2.0 nm) of SN 1998bw was obtained with the ESO 3.6-m telescope (+ EFOSC2) on Sept. 12.13 UT, 125 days after maximum light in B. The object has definitely entered its nebular phase, showing strong emissions of Na I D 589-nm, [O I] 630-nm, [Ca II] 728-nm, and the Ca II infrared triplet. Nevertheless, this supernova continues to be exceptional among the known type-Ic events: contrary to the expected spectral evolution in the blue region, SN 1998bw shows strong emissions centered at 400, 460, and 527 nm. While that at 400 nm is easily identified as Ca II H and K, the other two lines are puzzling; they are the most intense features after the [O I] 630-nm emission, and they have a FWHM of ∼ 16 nm. If they are due to Fe II (455.5, 521.5 nm), as the strong resemblance of this spectral region with that shown by type-Ia supernovae would suggest, we are looking at a very rare case of type-Iac supernova that shows emission features common to both type-Ia and type-Ib/c events. So far, only two other supernovae have exhibited such a strange behavior: SN 1990aj (see Piemonte 1997, “Observational properties of SNe Ib/c”,

http://athena.pd.astro.it/~supern/preprints.html) and SN 1993R

(Filippenko 1997, in Thermonuclear Supernovae, Canal et al., eds., Dordrecht: Kluwer, p. 795), but they suffer from scanty photometric data. We suggest a search for a possible coincidence of SN 1990aj or SN 1993R with known γ-ray bursts, to investigate further the claimed association of type-Ib/c supernovae with this class of events.”

1998 September 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green